Tax Shield:

1. The "tax shield" effect is best described as:

a. The fact that companies' NI is reduced by less than their interest expense.

b. The benefit companies get by relocating to low-tax jurisdictions.

c. The tax relief offered to many companies involved in Public/Private Partnership projects.

d. The effect of using legal shell companies to avoid taxes.

2.SloMo Films is considering taking out a $1,000K, 10% annual interest loan. It the company's tax rate is 25%, by how much will this loan reduce the company's annual Net Income?
Answer #1

rate positively ..

Ans 1 Correct answer is option :
a. The fact that companies' NI is reduced by less than their interest expense.
Ans 2 how much will this loan reduce the company's annual Net Income
i Annual interest = 100000
1000K*10%
ii=i*25% Tax shield @ 25% 25000
iii=i-ii Net reduction in net income = 75000